OpenStudy (anonymous):

Find the x-coordinates where f '(x) = 0 for f(x) = 2x + sin(2x) in the interval [0, 2π].

OpenStudy (freckles):

so you need to solve f'=0 in the interval [0,2pi] first find f'

OpenStudy (anonymous):

2cos(2x)+2 equals f'

OpenStudy (freckles):

\[2\cos(2x)+2=0 \\ 2\cos(2x)=-2 \\ \cos(2x)=-1\]

OpenStudy (freckles):

solve that equation on the interval 0 to 2pi

OpenStudy (anonymous):

i had this earlier but i could not figure out how to get rid of the 2x like it was

OpenStudy (freckles):

let u=2x so if x is between 0 and 2pi and x=u/2 then u is between 0 and 4pi so solve cos(u)=-1 on the interval [0,4pi]

OpenStudy (freckles):

cos(u)=-1 when u=pi can you find another value?

OpenStudy (anonymous):

If u equals pi then do you have to add the one and set it equal to zeros to solve?

OpenStudy (freckles):

i was asking you know another value between 0 and 4pi such that cos(u)=-1

OpenStudy (freckles):

I gave you one value for u that satisfies the equation

OpenStudy (freckles):

cos(u)=-1 gives us the solutions as u=pi or u=3pi in the interval [0,4pi]

OpenStudy (freckles):

since u=2x then we actually have to solve 2x=pi and solve 2x=3pi

OpenStudy (freckles):

when you solve those two equations for x you are done

OpenStudy (anonymous):

do i just do cos(2pi) and cos(2*3pi)?

OpenStudy (freckles):

no you just solve 2x=pi for x and solve 2x=3pi for x

OpenStudy (freckles):

divide both sides by 2

OpenStudy (anonymous):

the interval was 0 to 2pi, so would i just have pi

OpenStudy (freckles):

no the interval for x was 0 to 2pi but we let u=2x so that means we had u/2 is between 0 and 2pi so that means u is between 0 and 4pi so we solved cos(u)=-1 on the interval [0,4pi] giving us u=pi and u=3pi but we had u=2x so we have 2x=pi and 2x=3pi solving both of these equations gives you solutions for x only between 0 and 2pi

OpenStudy (freckles):

isn't x=pi/2 and x=3pi/2 both in the interval 0 to 2pi?

OpenStudy (freckles):

\[\cos(2x)=-1 ; 0 \le x \le 2 \pi \\ \\ \text{ if we let } u=2x \text{ then we have: } \\ \cos(u)=-1; 0 \le \frac{u}{2} \le 2\pi \\ \text{ but that inequality actually says } 0 \le u \le 4\pi (\text{ if you multiply both sides by 2}) \]